# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v3

@@ 5 Experiments
Qwen2.5-VL-3B Bai et al. (2025) Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
FFT with $D^{\text{VQA}}$ 75.3 (↑19.5%) 82.0 (↑0.2%) 74.7 (↑3.3%) 76.1 (↑5.0%)
@@
LLaVA-1.5-13B Liu et al. (2023a) Vanilla 65.8 (100%) 72.1 (100%) 64.5 (100%) 66.2 (100%)
   + KL 62.5 (↓5.0%) 74.6 (↑3.5%) 67.6 (↑4.8%) 68.0 (↑2.7%)
   + JS 65.8 (↑0.0%) 74.7 (↑3.6%) 67.6 (↑4.8%) 68.6 (↑3.6%)
   + RG 58.3 (↓11.4%) 73.5 (↑1.9%) 67.4 (↑4.5%) 66.9 (↑1.1%)
   + ADV 57.6 (↓12.5%) 74.2 (↑2.9%) 68.3 (↑5.9%) 67.5 (↑2.0%)
Ours 62.6 (↓4.9%) 73.7 (↑2.2%) 68.4 (↑6.0%) 68.4 (↑3.3%)
@@
Table 3: Evaluation of Caltech-101 (image-heavy) and MMLU (text-heavy) across different ablation study settings (full results in Table 6). Model Method Caltech-101 MMLU Orig Perturbed Orig Perturbed LLaVA-1.5-7B Vanilla 97.0 57.4 46.3 45.2 $D^{\text{VQA}}$ 96.2 46.3 46.8 45.6 $D^{\text{AUG}}$ 98.5 98.6 51.1 50.7    +KL 98.6 98.7 51.1 50.7    +ADV 98.7 98.5 50.6 50.3 Ours 99.3 99.0 51.5 51.0 InstructBLIP-7B Vanilla 90.3 17.5 35.3 35.2 $D^{\text{VQA}}$ 92.1 23.1 40.9 40.2 $D^{\text{AUG}}$ 99.0 56.1 50.0 49.7    +JS 98.9 98.4 50.7 49.3    +ADV 99.1 85.2 49.3 48.4 Ours 99.2 98.3 50.2 49.7 Table 4: OOD robustness evaluation on Caltech-101 and MMLU, with OCR noise and Screenshot distractors (full results in Table 11). Model Method Caltech-101 MMLU Orig OCR Orig Screenshot LLaVA-1.5-7B Vanilla 97.0 92.8 46.3 44.8 $D^{\text{AUG}}$ 98.5 98.0 51.1 50.6    +ADV 98.7 98.4 50.6 51.3 Ours 99.3 99.0 51.5 51.3 InstructBLIP-7B Vanilla 90.3 83.6 35.3 34.9 $D^{\text{AUG}}$ 99.0 98.4 50.0 49.2    +ADV 99.1 98.8 49.3 49.3 Ours 99.2 99.0 50.2 49.2 Qwen2.5-VL-7B Vanilla 99.1 99.2 69.3 57.0 $D^{\text{AUG}}$ 99.7 99.3 70.4 63.7    +ADV 99.6 99.6 70.4 69.7 Ours 99.6 99.5 69.8 69.7
@@
To evaluate the effectiveness of each component in our framework, we conduct a comprehensive ablation study across multiple models and scales. We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+RG (injecting random Gaussian noise into token embeddings), FFT+ADV (FFT with heuristic & adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).  Table 2 presents overall VQA performance, and  Table 3 evaluates model robustness under unimodal settings. ( Table 5 reports results with all perturbations.) Together, these results show the effectiveness of our method in improving both general VQA accuracy and robustness under modality interference. Across all model families (Qwen2.5-VL, InstructBLIP, LLaVA-1.5) and model sizes (3B/7B/13B), our method consistently achieves best overall performance, improving accuracy on both unimodal and multimodal benchmarks. For instance, it boosts overall VQA accuracy (e.g., +14.9% on InstructBLIP-7B), but also enhances robustness to modality interference–improve the performance under perturbations by over 50% on image-heavy tasks(e.g. 17.5% $\rightarrow$ 98.3% with InstructBLIP-7B on Caltech101). We also extend evaluation from MCQA to free-form QA (Section C.5).
@@
We observe consistent improvements across both unimodal and multimodal tasks when moving from FFT w/ $D^{\text{VQA}}$ to FFT w/ $D^{\text{AUG}}$, highlighting the importance of incorporating modality-specific supervision and heuristic perturbations. Building upon this, adding consistency regularization yields further gains by stabilizing model predictions under controlled perturbations on $X_{I}$ or $X_{T}$. Both KL and JS objectives lead to similar improvements, suggesting that the model equally benefits from all heuristic perturbations regardless of anchor choice.Finally, we compare adversarial perturbations with random Gaussian noise, and find that FFT+ADV consistently outperforms FFT+RG across most backbones, indicating that structured perturbations more effectively suppress spurious shortcuts and promote robust, task-relevant representations. To further validate generalization, we introduce two real-world out-of-distribution perturbations at test time: (i) noisy OCR snippets sampled from FUNSD (Jaume et al., 2019) as irrelevant text into image-heavy tasks; and (ii) unrelated UI screenshots from RICO (Deka et al., 2017) as distractor images in text-heavy tasks. As shown in Table 4, adversarial training significantly improves robustness under these unseen perturbations with consistent gains (e.g., on InstructBLIP-7B, 83.6% $\rightarrow$ 99.0% under OCR noise). These results demonstrate that the modest overhead of adversarial training (see Appendix E) yields substantial gains in out-of-domain generalization, a crucial property for reliable deployment.
@@ 6 Conclusion
Ethics Statement
Our research focuses on analyzing and mitigating modality interference in Multimodal Large Language Models. All experiments are conducted on publicly available benchmark datasets, including Mini-ImageNet, Caltech-101, OpenBookQA, MMLU, ScienceQA, MM-Bench, and Seed-Bench, which contain no personally identifiable or sensitive information beyond what is publicly released. We do not foresee direct risks of harm to individuals or groups arising from this research. Nevertheless, potential societal impacts include bias amplification or misinterpretation when deploying MLLMs in real-world applications. We note these risks and emphasize that our contributions are methodological and diagnostic rather than application-specific. No human subjects were involved, and no IRB approval was required. All funding sources are acknowledged in the main paper. Further discussions of limitations, broader impacts and LLM use are provided in Appendix Appendix F, Appendix Appendix G and Appendix Appendix H.
Reproducibility Statement
We are committed to ensuring the reproducibility of our work.
  • Code and Implementation: We will release a full open-source codebase, including data processing, training, and evaluation scripts, upon publication.
  • Datasets: All datasets used in this work are publicly available (Mini-ImageNet, Caltech-101, OpenBookQA, MMLU, ScienceQA, MM-Bench, Seed-Bench). Detailed preprocessing steps and dataset conversions into unified multiple-choice VQA format are described in  Section 5 and Appendix Appendix C.
  • Model and Training Details: Hyperparameters (learning rates, batch sizes, epochs, optimizer choices) and architectural specifications are reported in  Section 5 and Appendix Appendix B, Appendix D, Appendix E.
  • Evaluation: Metrics, baselines, and evaluation protocols are fully documented in  Section 5 and Appendix Appendix C with complete ablation results.
Together, these materials should enable independent researchers to reproduce our findings.
@@ Appendix A Appendix Summary
Detailed Experimental Results (Appendix C): This section includes three key tables— Table 5,  Table 6 and  Table 7—which report model performance on unimodal and multimodal tasks under various perturbation settings and ablation conditions(additional models included).  Table 8 records the performance of different vanilla MLLMs under modality interference across modality-heavy datasets. We also include radar plots ( Figure 4) that visualize task-wise robustness across different MLLMs. We provided the detailed experimental results on Qwen2.5-VL-7b (Bai et al., 2025) and InstructBlip-Vicuna-13b (Luo et al., 2023) and make further discussion on the selection of specific consistency loss. In  Table 11, we examine the generalization benefits of adversarial training by evaluating robustness under two types of out-of-distribution (OOD) perturbations: real-world OCR noise (from FUNSD (Jaume et al., 2019)) and unrelated screenshots (from RICO (Deka et al., 2017)). In  Table 9, we assess the impact of Chain-of-Thought prompting in mitigating modality interference, comparing its effectiveness against our method and standard baselines across both visual and textual modalities. In  Table 10, we report results on the free-form generative VQA benchmark TextVQA (Singh et al., 2019), highlighting our method’s generalizability beyond multiple-choice formats.
@@
Broader Impacts (Appendix G): We reflect on the ethical implications and societal benefits of our research. While our methods improve model robustness and alignment, we also acknowledge the dual-use nature of adversarial perturbations and advocate for safety-aware deployment.
LLM Use (Appendix H): Finally, we clarify that LLMs were only used to polish the writing of this paper.
@@
Qwen2.5-VL-3B Bai et al. (2025) Vanilla 98.9 94.9 98.8 94.4 79.9 74.6 63.5 61.1 72.5
FFT with $D^{\text{VQA}}$ 98.8 95.3 98.8 94.3 80.7 74.3 63.0 61.7 76.1
@@ Appendix C Detailed Experimental Results
Please see  Table 5,  Table 6,  Figure 4,  Table 8,  Table 7,  Table 11,  Table 9 and  Table 10 for more details.
@@
Qwen2.5-VL Models (Bai et al., 2025)
Qwen2.5-VL-3B Bai et al. (2025) Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
FFT with $D^{\text{VQA}}$ 75.3 (↑19.5%) 82.0 (↑0.2%) 74.7 (↑3.3%) 76.1 (↑5.0%)
@@
Ours 73.8 (↑17.1%) 81.5 (↓0.4%) 75.5 (↑4.4%) 76.4 (↑5.4%)
Qwen2.5-VL-7B Bai et al. (2025) Vanilla 85.1 (100%) 86.6 (100%) 77.0 (100%) 80.3 (100%)
@@
Note. While Qwen2.5-VL was originally instruction-tuned with proprietary in-house data (Bai et al., 2025), our reproduced version uses only publicly available LLaVA instruction-tuning data. Even under this constraint and without access to VQA-specific tuning samples, our models achieve comparable or even better performance across all VQA datasets—highlighting the robustness and effectiveness of our proposed perturbation-consistent fine-tuning strategy.
Table 8: Performance (%) of Vinilla models under modality interference across four datasets. We show accuracy under clean (origin) and various perturbations: Left: Mini-ImageNet and Caltech-101; Right: OpenBookQA and MMLU. UF = Unrelated Facts, MD = Misleading Descriptions, RP = Random Pixels, RI = Real Image, FB & FW = Full Black/White Canvas. (The results are averaged on multiple runs with standard deviation << 0.2)
@@
QwenVL2.5-7B (Bai et al., 2025) 99.3 99.3 96.5 99.3 99.4 96.8 86.2 80.9 86.3 86.4 69.5 67.6 69.1 69.1

### v4

@@ 4.3 Final Training Objective
Qwen2.5-VL-3B (Bai et al., 2025b) Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
FFT with $D^{\text{VQA}}$ 72.2 (↑14.6%) 82.0 (↑0.2%) 74.7 (↑3.3%) 75.5 (↑3.9%)
@@
InstructBLIP-13B (Luo et al., 2023) Vanilla 65.8 (100%) 72.1 (100%) 63.8 (100%) 65.8 (100%)
     + KL 67.2 (↑2.2%) 72.6 (↑0.7%) 64.0 (↑0.3%) 66.2 (↑0.6%)
     + ADV 66.1 (↑0.5%) 74.0 (↑2.6%) 64.1 (↑0.5%) 66.4 (↑0.9%)
Ours 66.2 (↑0.6%) 73.2 (↑1.5%) 64.3 (↑0.7%) 66.5 (↑1.1%)
5 Experiments
Models. We conduct experiments on three MLLM families with different parameter size: Qwen2.5-vl-3b (Bai et al., 2025b), LLaVA-1.5-7B & LLaVA-1.5-13B (Liu et al., 2023a) and InstructBLIP-Vicuna-7B (Luo et al., 2023). Following LLaVA and Qwen-VL, we freeze the vision encoder and train the multimodal projector and language model; for InstructBLIP, we instead freeze both the vision encoder and Q-Former, fine-tuning only the projection layer and language model (see Appendix B).
Baselines. We include following baselines for comparison: LLaVA-1.5-13B + I-MoF (Tong et al., 2024): By applying the designed Interleaved Mixture-of-Features (I-MoF) module on LLaVA-1.5-13B to spatially combine CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) visual tokens, it enhances visual grounding by integrating complementary features from contrastive and self-supervised vision encoders. VLMClassifier (Zhang et al., 2024): it enhances visually-grounded language models for image classification by fine-tuning them on ImageNet (Deng et al., 2009) (VLMClassifier-1) or ImageNet combining LLaVA-Instruct (Liu et al., 2023b) (VLMClassifier-2). R1-OneVision-7B (Yang et al., 2025): a reasoning-enhanced “thinking” model built upon the Qwen2.5-VL-7B backbone. We include it to investigate whether intrinsic reasoning capabilities can inherently mitigate modality interference. Chain-of-Thought (CoT) Prompting (Wei et al., 2022): we further evaluate prompt-based mitigation by encouraging structured reasoning through CoT-style prompting. The specific prompt design and results are reported in Section C.4.
Datasets. We evaluate models on benchmarks covering three task types: (i) Image-heavy tasks: Mini-ImageNet (Russakovsky et al., 2015) and Caltech-101 (Fei-Fei et al., 2004), used in both training and evaluation; (ii) Text-heavy tasks: OpenBookQA (Mihaylov et al., 2018) and MMLU (Hendrycks et al., 2020), used in both training and evaluation; (iii) VQA tasks: For training, we use LLaVA-Instruct-dataset (Liu et al., 2023b) as the instruction-tuning dataset for related models. For InstructBLIP, we additionally use TextCaps (Sidorov et al., 2020) following its original training configuration. For Qwen2.5-VL, whose instruction-tuning data is proprietary, we adopt LLaVA-Instruct as a standardized alternative. For evaluation, we adopt three multiple-choice VQA benchmarks: ScienceQA-IMG (Lu et al., 2022), MM-Bench-EN (Liu et al., 2023c), and Seed-Bench-IMG (Li et al., 2023). For ScienceQA and Seed-Bench, we only include examples with image context. For MM-Bench, we use the English version. All datasets are converted into a unified multiple-choice VQA format. We report the accuracy of model predictions and quantify the causal effect with the prediction change rate $\delta_{\text{cp}}$. All models are fine-tuned for 1 epoch with a fixed batch size $N_{\text{batch}}$. All the hyperparameters are listed in Appendix D. For each dataset, results are averaged over multiple independent runs. Following standard practice, inference is performed with deterministic decoding (temperature fixed at 0).
Achieving Pareto-Optimality Across Unimodal and Multimodal Tasks
As shown in Table 1, our method outperforms all baselines across different base MLLMs, demonstrating stronger robustness to modality interference. While CoT slightly improves performance in certain VQA settings, its overall gains are minimal and inconsistent, and it fails to mitigate modality interference (e.g., 85.9% vs. 97.9% on Mini-ImageNet). Similarly, even the reasoning-enhanced R1-OneVision-7B suffers from significant degradation (e.g., dropping from 96.5% to 84.7% on Caltech-101), indicating that intrinsic “thinking” capabilities alone are insufficient to resolve the issue. While I-MoF enhances visual grounding by integrating multiple visual features, it still suffers from modality interference: e.g. LLaVA-1.5-13B + I-MoF drops from 93.9% to 70.1% under perturbation (↓23.8%), indicating reliance on spurious textual cues. In contrast, our method maintains perturbed performance at 98.4% (↓0.1%). On the other hand, VLMClassifier, adopts vision-only fine-tuning, which leads to critical limitations–vulnerability to cross-modal interference and degradation on VQA tasks, as LLaVA-1.5-7b + VLMClassifier-1 only reaches 35.8%/36.2% on MM-Bench/SeedBench, notably lower than both base LLaVA and our method (73.7%/68.4%), highlighting that vision-centric strategies, without addressing modality alignment, are insufficient for robust multimodal understanding. In text-heavy tasks, our method also achieves superior perturbed performance, indicating that addressing modality interference directly, rather than merely improving representations, is key to robust multimodal reasoning. Overall, unlike prior methods that often trade off between unimodal and multimodal performance, our method consistently improves both, achieving Pareto-optimality.
@@ 5 Experiments
Table 3: Evaluation of Caltech-101 (image-heavy) and MMLU (text-heavy) across different ablation study settings across different model families (full results in Table 6).
Model Method Caltech-101 MMLU
Orig Perturbed Orig Perturbed
LLaVA-1.5-7B Vanilla 97.0 57.4 62.4 56.4
$D^{\text{VQA}}$ 96.2 46.3 61.3 55.5
$D^{\text{AUG}}$ 98.5 98.6 78.6 77.2
   +KL 98.6 98.7 81.4 81.2
   +ADV 98.7 98.5 81.7 80.8
Ours 99.3 99.0 81.8 81.5
InstructBLIP-7B Vanilla 90.3 17.5 50.9 46.2
$D^{\text{VQA}}$ 92.1 23.1 49.8 45.2
$D^{\text{AUG}}$ 99.0 56.1 75.0 74.8
   +JS 98.9 98.4 78.0 76.6
   +ADV 99.1 85.2 76.8 76.3
Ours 99.2 98.3 79.0 78.3
Table 4: Evaluation of out-of-distribution (OOD) robustness under unseen perturbation types across both image-heavy and text-heavy tasks on all evaluated datasets (Seen: Mini-ImageNet, Caltech-101, OpenBookQA, MMLU; Unseen: Food-101, ARC Challenge).
Model Method Image-heavy Text-heavy
Mini-ImageNet Caltech-101 Food-101 OpenBookQA MMLU ARC Challenge
Orig OCR Orig OCR Orig OCR Orig Screenshot Orig Screenshot Orig Screenshot
LLaVA-1.5-7B Vanilla 95.3 88.9 97.0 92.8 90.2 83.6 62.4 50.2 46.3 44.8 53.4 51.8
$D^{\text{AUG}}$ 98.2 98.0 98.5 98.0 91.3 83.5 78.6 78.0 51.1 50.6 55.7 53.8
   +ADV 98.7 98.2 98.7 98.4 91.8 86.7 81.7 81.3 50.6 51.3 57.9 57.4
Ours 98.6 98.2 99.3 99.0 92.1 89.5 81.8 81.2 51.5 51.3 62.8 62.6
InstructBLIP-7B Vanilla 92.0 81.4 90.3 83.6 67.9 48.0 50.9 40.0 35.3 34.9 33.7 36.5
$D^{\text{AUG}}$ 98.5 95.7 99.0 98.4 55.5 56.5 75.0 74.4 50.0 49.2 27.3 30.7
   +ADV 98.7 97.5 99.5 98.8 93.2 85.9 76.8 76.2 49.3 49.3 50.1 44.5
Ours 98.4 97.3 99.2 99.0 95.6 88.1 79.0 77.8 50.2 49.2 57.6 57.0
Qwen2.5-VL-7B Vanilla 99.3 99.2 99.1 99.2 97.1 96.7 85.9 69.3 69.3 57.0 85.6 79.6
$D^{\text{AUG}}$ 99.6 99.5 99.7 99.3 97.1 96.5 90.2 85.8 70.4 63.7 82.9 80.2
   +ADV 99.4 99.5 99.6 99.6 97.3 96.9 91.7 91.5 70.4 69.7 85.0 81.2
Ours 99.6 99.5 99.6 99.5 97.5 97.4 90.9 90.7 69.8 69.7 84.1 82.6
@@
To evaluate the effectiveness of each component in our framework, we conduct a comprehensive ablation study across multiple models and scales. We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+ADV (FFT with heuristic & adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).  Table 2 presents overall VQA performance, and  Table 3 evaluates model robustness under unimodal settings. ( Table 5 reports results with all perturbations.) Together, these results show the effectiveness of our method in improving both general VQA accuracy and robustness under modality interference. Across all model families (Qwen2.5-VL, InstructBLIP, LLaVA-1.5) and model sizes (3B/7B/13B), our method consistently achieves best overall performance, improving accuracy on both unimodal and multimodal benchmarks. For instance, it boosts overall VQA accuracy (e.g., +14.9% on InstructBLIP-7B), but also enhances robustness to modality interference–improve the performance under perturbations by over 50% on image-heavy tasks(e.g. 17.5% $\rightarrow$ 98.3% with InstructBLIP-7B on Caltech101). We also extend evaluation from MCQA to free-form QA and additional diagnostic experiments, including a no-mask ablation, which exhibit consistent robustness trends (see Sections C.5 and C.7).
@@
We observe consistent improvements across both unimodal and multimodal tasks when moving from FFT w/ $D^{\text{VQA}}$ to FFT w/ $D^{\text{AUG}}$, highlighting the importance of incorporating modality-specific supervision and heuristic perturbations. Building upon this, adding consistency regularization yields further gains by stabilizing model predictions under controlled perturbations on $X_{I}$ or $X_{T}$. Both KL and JS objectives lead to similar improvements, suggesting that the model equally benefits from all heuristic perturbations regardless of anchor choice. To further validate generalization, we introduce two real-world out-of-distribution perturbations at test time: (i) noisy OCR snippets sampled from FUNSD (Jaume et al., 2019) as irrelevant text into image-heavy tasks; and (ii) unrelated UI screenshots from RICO (Deka et al., 2017) as distractor images in text-heavy tasks. Beyond existing datasets, we further evaluate on Food-101 (Bossard et al., 2014) (unseen image-heavy dataset) and ARC-Challenge (Clark et al., 2018) (unseen text-heavy dataset), neither of which is used during training. As shown in Table 4, our method significantly improves robustness under these unseen perturbations with consistent gains (e.g., on InstructBLIP-7B, 83.6% $\rightarrow$ 99.0% under OCR noise on Caltech-101). Crucially, results on new datasets confirm our method’s strong OOD generalization. Under misleading text (seen perturbation) and OCR-style noise (unseen perturbation), vanilla models suffer severe degradation, while our method preserves accuracy close to clean-baseline performance. In contrast, heuristic data augmentation alone fails to generalize to unseen noise, and may even degrade performance (e.g., on InstructBLIP-7B, heuristic augmentation degrades performance on Food-101 under clean inputs (67.9% vs. 55.5%). These results demonstrate that the modest overhead of adversarial training (see Appendix E) yields substantial gains in out-of-domain generalization (full results in Table 11 and Table 12).
@@ 6 Conclusion
7 Impact Statement
This paper presents work whose goal is to advance the field of machine learning by improving the understanding and robustness of multimodal large language models. The proposed analyses and training methods aim to make model behavior more reliable under modality-specific perturbations.
The societal impacts of this work are consistent with those commonly associated with advances in machine learning research, and we do not anticipate any unique ethical concerns that require special discussion.
@@ Appendix A Appendix Summary
Detailed Experimental Results (Appendix C): This section includes three key tables— Table 5,  Table 6 and  Table 7—which report model performance on unimodal and multimodal tasks under various perturbation settings and ablation conditions(additional models included).  Table 8 records the performance of different vanilla MLLMs under modality interference across modality-heavy datasets. We also include radar plots ( Figure 5) that visualize task-wise robustness across different MLLMs. We provided the detailed experimental results on Qwen2.5-VL-7b (Bai et al., 2025b) and InstructBLIP-Vicuna-13b (Luo et al., 2023) and make further discussion on the selection of specific consistency loss. In  Table 12,  Table 11, we examine the generalization benefits of adversarial training by evaluating robustness under two types of out-of-distribution (OOD) perturbations: real-world OCR noise (from FUNSD (Jaume et al., 2019)) and unrelated screenshots (from RICO (Deka et al., 2017)) and two new datasets: Food-101 (Bossard et al., 2014) and ARC Challenge (Clark et al., 2018). In  Table 9, we assess the impact of Chain-of-Thought prompting in mitigating modality interference, comparing its effectiveness against our method and standard baselines across both visual and textual modalities. In  Table 10, we report results on the free-form generative VQA benchmark TextVQA (Singh et al., 2019), highlighting our method’s generalizability beyond multiple-choice formats. In  Table 13, we make another ablation study without modality-specific masking.
@@
LLM Use (Appendix G): Finally, we clarify that LLMs were only used to polish the writing of this paper.
@@
Qwen2.5-VL-3B (Bai et al., 2025b) Vanilla 98.9 94.9 98.8 94.4 79.9 74.6 63.5 61.1 72.5
FFT with $D^{\text{VQA}}$ 98.8 95.3 98.8 94.3 80.7 74.3 63.0 61.7 75.5
@@ Appendix C Detailed Experimental Results
Please see  Table 5,  Table 6,  Figure 5,  Table 8,  Table 7,  Table 12,  Table 11,  Table 9 and  Table 10 for more details. In table  Table 7: We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+RG (injecting random Gaussian noise into token embeddings), FFT+ADV (FFT with heuristic & adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).
@@
Qwen2.5-VL Models (Bai et al., 2025b)
Qwen2.5-VL-3B (Bai et al., 2025b) Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
FFT with $D^{\text{VQA}}$ 72.2 (↑14.6%) 82.0 (↑0.2%) 74.7 (↑3.3%) 75.5 (↑3.9%)
@@
Ours 73.8 (↑17.1%) 82.1 (↑0.1 %) 75.5 (↑4.4%) 76.4 (↑5.4%)
Qwen2.5-VL-7B (Bai et al., 2025b) Vanilla 85.1 (100%) 86.6 (100%) 77.0 (100%) 80.3 (100%)
@@
Note. While Qwen2.5-VL was originally instruction-tuned with proprietary in-house data (Bai et al., 2025b), our reproduced version uses only publicly available LLaVA instruction-tuning data. Even under this constraint and without access to VQA-specific tuning samples, our models achieve comparable or even better performance across all VQA datasets—highlighting the robustness and effectiveness of our proposed perturbation-consistent fine-tuning strategy.
Table 8: Performance (%) of V a lin na models under modality interference across four datasets. We show accuracy under clean (origin) and various perturbations: Left: Mini-ImageNet and Caltech-101; Right: OpenBookQA and MMLU. UF = Unrelated Facts, MD = Misleading Descriptions, RP = Random Pixels, RI = Real Image, FB & FW = Full Black/White Canvas. (The results are averaged on multiple runs with standard deviation << 0.2)
@@
QwenVL2.5-7B (Bai et al., 2025b) 99.3 99.3 96.5 99.3 99.4 96.8 86.2 80.9 86.3 86.4 69.5 67.6 69.1 69.1
QwenVL3-8B (Bai et al., 2025a) 99.3 98.8 97.2 99.8 99.7 98.1 87.0 85.7 87.5 88.1 75.6 75.0 76.2 76.3
Figure 4: Performance degradation under irrelevant perturbations reveals modality interference in MLLMs. Left: Mini-ImageNet (image-heavy) with Original input, Unrelated Facts, and Misleading Descriptions. Right: OpenBookQA (text-heavy) with Random Pixels, Full Black Canvas, and Irrelevant Real Images. Misleading descriptions most severely affect image-heavy tasks, while irrelevant real images cause the largest drop in text-heavy reasoning.